### v2

# softmax is not enough (for sharp out-of-distribution)

Abstract
A key property of reasoning systems is the ability to make sharp decisions on their input data. For contemporary AI systems, a key carrier of sharp behaviour is the softmax function, with its capability to perform differentiable query-key lookups. It is a common belief that the predictive power of networks leveraging softmax arises from “circuits” which sharply perform certain kinds of computations consistently across many diverse inputs. However, for these circuits to be robust, they would need to generalise well to arbitrary valid inputs. In this paper, we dispel this myth: even for tasks as simple as finding the maximum key, any learned circuitry must disperse as the number of items grows at test time. We attribute this to a fundamental limitation of the softmax function to robustly approximate sharp functions, prove this phenomenon theoretically, and propose adaptive temperature as an ad-hoc technique for improving the sharpness of softmax at inference time.
1 Motivation
It is no understatement to say that the $\mathtt{softmax}_{\theta}:\mathbb{R}^{n}\rightarrow[0,1]^{n}$ function [^1] Strictly speaking, the proper name for this function should be softargmax. We choose to retain the terminology introduced by [Bri89], primarily for reasons of alignment with modern deep learning frameworks.:
$\mathtt{softmax}_{\theta}(\mathbf{e})=\begin{bmatrix}\dfrac{\exp(e_{1}/\theta)}{\sum_{k}\exp(e_{k}/\theta)}&\dots&\dfrac{\exp(e_{n}/\theta)}{\sum_{k}\exp(e_{k}/\theta)}\end{bmatrix}$ (1)
is one of the most fundamental functions in contemporary artificial intelligence systems.
The role of softmax in deep learning is to convert any vector of logits, $\mathbf{e}\in\mathbb{R}^{n}$, into a probability distribution, in a form that is part of the exponential family. Further, softmax allows for application of a temperature parameter, $\theta\in\mathbb{R}$, to adjust the amount of probability mass attached to the highest logit—a concept borrowed from the Boltzmann distribution in statistical mechanics.
Initially, the primary utilisation of softmax in deep learning was within the final layer of classifiers. Its influence in this domain vastly expanded after it saw use in the internal layers—as a differentiable key-value store [GWD14] or a mechanism for attending over the most relevant parts of the input [BCB15]. This attentional framing of softmax was critical in defining important models for sequences [VSP+17, Transformers], images [DBK+21, ViTs] and graphs [VCC+18, GATs].
Several efforts attribute the success of softmax to its capability of modelling computations relevant to reasoning. This can be related to the concept of circuits in theoretical computer science [AB09]. Several interpretable pieces of “circuitry” [OCS+20] have already been discovered in large Transformers, primarily under the umbrella of mechanistic interpretability [ENO+21, OEN+22, WVC+22].
Here we study the robustness of such circuitry, especially when going beyond the distribution the models are trained on—a critical regime for reasoning engines. We find that, in spite of its many successes, softmax does not have a chance to robustly generalise such circuits out of distribution, especially because it provably cannot approximate sharpness with increasing problem size (Figure 1). Here we call a function sharp if its output only depends on a constant number of its inputs (e.g. max).
Figure 1: Illustration of Theorem 2.2, one of our key results. Assuming a tokenised input from a fixed vocabulary and a non-zero temperature, for every softmax attention head inside an architecture comprising only MLPs and softmax self-attention layers, it must hold that, given sufficiently many tokens, its attention coefficients will disperse, even if they were sharp for in-distribution instances.
We hope that this result encourages future study of alternative attentional functions, in light of the problems we identify, especially for building reasoning engines of the future. That being said, we also believe that our findings indicate ways to modify the softmax function to support sharpness for longer—as one simple example of this, we propose an adaptive temperature mechanism for softmax.
Background
The analysis of attentional coefficients and attempting to attribute interpretable operations to them dates back to the earliest deployments of internal softmax layers at scale; examples include [GWD14, Figure 6], [BCB15, Figure 3], [VSP+17, Figures 3–5] and [QTM+18, Figure 5]. A strong current in this space analyses the self-attentional heads of Transformers [VTM+19, JW19].
With the rise of large language models, mechanistic interpretability has taken charge in detecting and elucidating various circuits in Transformers [ENO+21]. Some prominent discoveries include induction heads [OEN+22], indirect object identification [WVC+22], multiple-choice heads [LRK+23], successor heads [GOOC23], attentional sinks [DOMB23], comparator heads [HLV24] and retrieval heads [WWX+24]. Most recently, these efforts have relied on sparse autoencoders [KKB+24].
While the skills above are quite impressive and span many rules one might hope a robust reasoning system would have, and the discovered heads always appear sharp when inspected, it is also known that many easy tasks requiring sharp attention—such as finding minima—are hard to do reliably with LLMs out-of-distribution [MMI+24, Figure 6]. More challenging sharp order statistic tasks, such as finding the second minimum [OV22] may even be hard to learn in-distribution. The discrepancy of such results with the previous paragraph motivate our study, and formalisation of softmax dispersion.
Certain dispersion effects in softmax—e.g. as an effect of increasing temperature—are already well-understood in thermodynamics. A core contribution of our work is understanding dispersion in a setting where the amount of logits can vary, which is relevant for generalisation in Transformers. We are not the first to observe dispersion in this setting empirically; prior works studying the capability of Transformers to execute algorithms [YSK+20] and perform random-access lookups [EPM24] also note dispersion patterns. Our work is the first to rigorously prove these effects, directly attribute them to the softmax operator, as well as propose ways to improve sharpness empirically within softmax. The proof technique we will use to demonstrate this is inspired by [BBK+24], though unlike their work, our results apply regardless of whether the computational graph is bottlenecked or not.
2 Demonstrating and proving the dispersion in softmax and Transformers
To motivate our theory, we train a simple architecture including a single dot-product attention head to predict a feature of the maximum item in a set. Each item’s features are processed with a deep MLP before attending, and the output vector of the attention is passed to a deep MLP predictor (see Appendix A for experimental details). We train this model using sets of $\leq 16$ items, and in Figure 2 we visualise the head’s attentional coefficients, computed over sets of varying size at inference time.
Figure 2: Visualising the attentional head for the max retrieval task for a batch of $32$ sets, over the $16$ items with largest key. If the head operates correctly, it must allocate sharp attention to the rightmost item. From left to right, in each frame we double the number of items the head has to process.
Figure 3: Entropy of attention heads in the first block of Gemma 2B with prompt ‘‘What is the maximum in the following sequence: {seq}? The maximum is:’’ and varying the number of elements in seq.
While the model indeed attributes focus sharply and cleanly on the maximum item, this only holds true on the problem sizes that the model was trained on. As we simulate an out-of-distribution setting where the problem size increases (without changing the value distribution), the attentional coefficients eventually disperse towards the uniform distribution.
This effects manifests in the attention heads of Transformers as well—we visualise the entropy (a proxy for sharpness) of Gemma 2B [GMH+24]’s heads when answering a similar maximisation task in Figure 3.
In fact, we can show that this effect is inevitable in softmax using the following Lemma, which we prove below:
Lemma 2.1 (softmax must disperse).
Let $\mathbf{e}^{(n)}\in\mathbb{R}^{n}$ be a collection of $n$ logits going into the $\mathtt{softmax}_{\theta}$ function with temperature $\theta>0$, bounded above and below s.t. $m\leq e^{(n)}_{k}\leq M$ for some $m,M\in\mathbb{R}$. Then, as more items are added ($n\rightarrow+\infty$), it must hold that, for each item $1\leq k\leq n$, $\mathtt{softmax}_{\theta}(\mathbf{e}^{(n)})_{k}=\Theta(\frac{1}{n})$. That is, the computed attention coefficients disperse for all items.
Proof.
Let us denote the attentional coefficient assigned to $k$ by $\alpha^{(n)}_{k}=\texttt{softmax}_{\theta}(\mathbf{e}^{(n)})_{k}\in[0,1]$. Then we can bound $\alpha^{(n)}_{k}$ above as:
$\alpha^{(n)}_{k}=\frac{\exp(e^{(n)}_{k}/\theta)}{\sum_{l}\exp(e^{(n)}_{l}/\theta)}\leq\frac{\exp(M/\theta)}{n\exp(m/\theta)}=\frac{1}{n}\exp\left(\frac{M-m}{\theta}\right)$ (2)
Similarly, we can bound $\alpha^{(n)}_{k}$ below as:
$\alpha^{(n)}_{k}=\frac{\exp(e^{(n)}_{k}/\theta)}{\sum_{l}\exp(e^{(n)}_{l}/\theta)}\geq\frac{\exp(m/\theta)}{n\exp(M/\theta)}=\frac{1}{n}\exp\left(\frac{m-M}{\theta}\right)$ (3)
Hence, if we let $\delta=(M-m)$
$\frac{1}{n}\exp-\frac{\delta}{\theta}\leq\alpha^{(n)}_{k}\leq\frac{1}{n}\exp\frac{\delta}{\theta}$ (4)
Which implies $\alpha_{k}^{(n)}=\Theta(\frac{1}{n})$ as $\delta$ and $\theta$ are both constants. ∎
Lemma 2.1 relies on being able to bound the logit values with specific constants. In modern Transformer architectures operating over a vocabulary of possible token values, we can actually bound the logits in every single attentional layer—implying that dispersion must happen everywhere in a Transformer for sufficient problem sizes. We prove this important result now:
Theorem 2.2 (softmax in Transformers over vocabularies must disperse).
Let $\mathcal{X}\subset\mathbb{R}^{m}$ be an $m$-dimensional input feature space, and let $\mathbf{X}^{(n)}\in\mathcal{X}^{n}$ be a matrix of input features for $n$ items. Further, assume that input features come from a finite set of possible values, i.e. $|\mathcal{X}|<|\mathbb{N}|$. Let $e_{j}^{(n)}=(\mathbf{q}^{(n)})^{\top}\mathbf{k}_{j}^{(n)}$ where $\mathbf{q}^{(n)}=\phi(\mathbf{x}^{(n)}_{1},\dots,\mathbf{x}^{(n)}_{n})$ and $\mathbf{K}^{(n)}=\kappa(\mathbf{x}^{(n)}_{1},\dots,\mathbf{x}^{(n)}_{n})$, where $\phi:\mathcal{X}^{n}\rightarrow\mathbb{R}^{k}$ and $\kappa:\mathcal{X}^{n}\rightarrow\mathbb{R}^{n\times k}$ are continuous functions, each expressible as a composition of $L$ layers $g_{L}\circ f_{L}\circ\dots\circ g_{1}\circ f_{1}$ where each layer contains a feedforward component $f_{i}(\mathbf{z}_{1},\dots,\mathbf{z}_{n})_{k}=f_{i}(\mathbf{z}_{k})$ or a self-attentional component $g_{i}(\mathbf{z}_{1},\dots,\mathbf{z}_{n})_{k}=\sum_{1\leq l\leq n}\alpha_{lk}v_{i}(\mathbf{z}_{l})$ where $\alpha_{lk}\in[0,1]$ are softmax-normalised attention coefficients and $v_{i}$ is a feedforward network. Then, for any $\theta>0$ and $\epsilon>0$, there must exist an $n\in\mathbb{N}$ such that $\mathtt{softmax}_{\theta}(\mathbf{e}^{(n)})_{k}<\epsilon$ for all $1\leq k\leq n$. That is, attention coefficients must disperse in all Transformer heads if the input vocabulary is finite.
Proof.
Firstly, note that since $\mathcal{X}$ is a finite set of $m$-dimensional vectors, then it is also part of a compact space spanning all convex combinations of those vectors. Then, all feedforward layers, $f_{i}$ and $v_{i}$, being continuous functions, move inputs from a compact set to another compact set. Similarly, every self-attentional layer, $g_{i}$, computes a convex combination of the outputs of $v_{i}$, and as such, if outputs of $v_{i}$ are on a compact space, the outputs of $g_{i}$ remain on the same compact space. Therefore, if the input space of $\phi$ and $\kappa$ is compact, then the output space of $\phi$ and (each row of) $\kappa$ on $\mathbb{R}^{k}$ must be compact as well, regardless of the choice of $n$. Further, the dot product of two vectors $(\mathbf{q}^{(n)})^{\top}\mathbf{k}_{j}^{(n)}$ coming from compact spaces must be compact as well. Hence, by definition, the logits must be bounded by $m\leq e^{(n)}_{k}\leq M$ for constant $m$ and $M$. Then, letting $\delta=M-m$, we know (Lemma 2.1) that $\mathtt{softmax}_{\theta}(\mathbf{e}^{(n)})_{k}\leq\frac{1}{n}\exp\left(\delta/\theta\right)$, so for all $n>\dfrac{\exp\left(\delta/\theta\right)}{\epsilon}$ this value will be below $\epsilon$. ∎
3 Adaptive temperature
Since we now know dispersion is inevitable, are there any ways we can leverage our theory’s findings to make softmax sharper? One obvious constraint our theory rests on is the assumption that $\theta>0$, i.e. that our temperature is nonzero. While zero temperature—also known as hard attention [DBLdF12, Ran14, MHG+14, XBK+15]—guarantees sharpness, training large-scale Transformers with it tends to not work well in practice [BIB+24].
What about applying zero temperature to an already-trained Transformer? We can show that this is also problematic since, for any attention head where the Transformer has learnt to induce sharpness, it necessarily did so by increasing magnitude of its weights (see Appendix B for a proof):
Proposition 3.1 (Sharpness in Transformers necessitates large weights).
Let $\mathbf{e}^{(n)}\in\mathbb{R}^{n}$ be a collection of $n$ logits, computed using a dot product attention mechanism; i.e. $e^{(n)}_{k}=\left\langle\mathbf{Q}\mathbf{y},\mathbf{K}\mathbf{x}_{k}\right\rangle$, where $\mathbf{y}\in\mathbb{R}^{m}$ is a query vector and $\mathbf{Q},\mathbf{K}\in\mathbb{R}^{m^{\prime}\times m}$ are parameters. Let $\delta=\max\limits_{1\leq i\leq n}e^{(n)}_{i}-\min\limits_{1\leq j\leq n}e^{(n)}_{j}$ be their maximum difference. Then $\delta$ is upper bounded as $\delta\leq 2\sigma_{\mathrm{max}}^{(Q)}\sigma_{\mathrm{max}}^{(K)}\|\mathbf{y}\|\max_{1\leq i\leq n}\|\mathbf{x}_{i}\|$, where $\sigma_{\mathrm{max}}^{(Q)},\sigma_{\mathrm{max}}^{(K)}\in\mathbb{R}$ are the largest singular values of $\mathbf{Q}$ and $\mathbf{K}$. That is, the sharpness of the softmax in Transformers depends on the norm of its parameters.
Note that there is a common practice of leveraging operators such as layer normalisation [BKH16] extensively within Transformer architectures, which clamps $\|\mathbf{x}_{i}\|$ and $\|\mathbf{y}\|$ if applied right before the query-key mechanism, more directly accentuating the impact of the singular values of $\mathbf{Q}$ and $\mathbf{K}$.
However, forcing large parameters promotes overfitting, and the likelihood that the incorrect item gets the largest logit—see Figure 2. Setting temperature to zero will then degrade accuracy—we might prefer to make the coefficients sharper while making sure that the chosen item is not left behind. This motivates our use of adaptive temperature, where we vary $\theta$ depending on the entropy in the input coefficients. Adaptive temperature can be elegantly motivated by the fact that decreasing the temperature must monotonically decrease the entropy, which is well-known in thermodynamics:
Proposition 3.2 (Decreasing temperature decreases entropy).
Let $\mathbf{e}^{(n)}\in\mathbb{R}^{n}$ be a collection of $n$ logits. Consider the Boltzmann distribution over these $n$ items, $p_{i}\propto\exp(-\beta e_{i}^{(n)})$ for $\beta\in\mathbb{R}$, and let $H=-\sum_{i}p_{i}\log p_{i}$ be its Shannon entropy. Then, as $\beta$’s magnitude increases, $H$ must monotonically decrease. Thus, since $\beta\propto\frac{1}{\theta}$ where $\theta$ is the temperature in $\mathtt{softmax}_{\theta}$, decreasing the temperature must monotonically decrease the entropy.
Figure 4: Entropy of the $\mathtt{softmax}_{\theta}$ function for 10 elements of a power series. Entropy increases with temperature but the rate at which it increases is heavily dependent on the attention logit distribution. N.B. degenerate cases: near $\lambda=0$ and $\lambda=1$ all logits are the same, leading to highest entropy.
Figure 5: The polynomial fit used to derive our adaptive formula for $\theta$ as a function of the Shannon entropy, $H$. The fit degree-4 function was $\theta\approx 1/(-1.791+4.917H-2.3H^{2}+0.481H^{3}-0.037H^{4})$. We do not apply the correction to $\theta$ if predicted greater than $1$.
We provide a full proof in Appendix C. To supplement Proposition 3.2 empirically, we also provide—in Figure 4—a visualisation of how the Shannon entropy varies with temperature, for a $10$-logit input with varying spread between the logits.
To compute the approximate temperature value as a function of entropy, we generate a dataset of inputs to our model where the maximal items do not obtain the highest logit. For each such input, we find the “optimal” value of $\theta$ that would maximise its probability. Then we fit an inverse degree-4 polynomial to this data—see Figure 5—and use it to predict temperatures to use at inference time. Note we do not wish to increase entropy; as such, we do not correct $\theta$ to values greater than $1$.
The JAX [BFH+18] implementation of our adaptive-$\theta$ softmax is provided below, and we use it as a drop-in replacement for jax.nn.softmax in all of our experiments. Note that it requires two calls to jax.nn.softmax in place of one.
{Verbatim}
[commandchars=
{}] def adaptive_temperature_softmax(logits): original_probs = jax.nn.softmax(logits)
poly_fit = jnp.array([-0.037, 0.481, -2.3, 4.917, -1.791]) # see Figure 5 entropy = jnp.sum(-original_probs * jnp.log(original_probs + 1e-9), axis=-1, keepdims=True) # compute the Shannon entropy beta = jnp.where( # beta = 1 / theta entropy > 0.5, # don’t overcorrect low-entropy heads jnp.maximum(jnp.polyval(poly_fit, entropy), 1.0), # never increase entropy 1.0)
return jax.nn.softmax(logits * beta)
Note that we are not the first to propose dynamically adapting temperature—[NZV18, RKH+21] do this in the classification layer (and hence do not have to handle an ever-increasing amount of items), whereas [CC22] perform it over intermediate attentional heads, but in a way that only depends on problem size (multiplying logits by $\log n$) and hence not taking into account initial logit sharpness.
4 Experimental results
To validate the utility of our proposed adaptive temperature scheme, we evaluate it on both our previously-mentioned max retrieval task—which allows us a pristine environment for evaluating whether adaptive temperature leads to more useful attention heads—as well and the CLRS-Text algorithmic reasoning benchmark for language models [MMI+24], which represents a challenging reasoning task for decoder-only Transformers, and is hence likely to require low-entropy behaviour.
Table 1: Improvements observed when applying adaptive temperature on the max retrieval task (without changing the parameters), averaged over ten seeds. $p$-values computed using a paired $t$-test.
ID size Out-of-distribution sizes
Model $16$ $32$ $64$ $128$ $256$ $512$ $1,024$ $2,048$ $4,096$ $8,192$ $16,384$
Baseline $\mathbf{98.6}\%$ $\mathbf{97.1}\%$ ${94.3}\%$ $89.7\%$ $81.3\%$ $70.1\%$ $53.8\%$ $35.7\%$ $22.6\%$ $15.7\%$ $12.4\%$
Adaptive $\theta$ $\mathbf{98.6}\%$ $\mathbf{97.1}\%$ $\mathbf{94.5}\%$ $\mathbf{89.9}\%$ $\mathbf{82.1}\%$ $\mathbf{72.5}\%$ $\mathbf{57.7}\%$ $\mathbf{39.4}\%$ $\mathbf{24.9}\%$ $\mathbf{17.5}\%$ $\mathbf{14.0}\%$
$p$-value $0.4$ $0.4$ $\underline{0.002}$ $\underline{2\cdot 10^{-5}}$ $\underline{2\cdot 10^{-4}}$ $\underline{3\cdot 10^{-5}}$ $\underline{10^{-4}}$ $\underline{6\cdot 10^{-4}}$ $\underline{0.02}$ $\underline{10^{-3}}$ $\underline{4\cdot 10^{-3}}$
4.1 max retrieval
For this task, we first train our single attention head architecture as described in Appendix A; then, we evaluate it at various numbers of input items, with and without applying adaptive temperature to its sole softmax function call. Note that this is a “pure” inference time adjustment—no modifications to the learned parameters are performed!
The results—averaged over ten seeds and with statistical significance tests applied—are summarised in Table 1. As is evident, applying adaptive temperature leads to a more performant retrieval head on out-of-distribution inputs, with statistical significance ascertained via a paired $t$-test.
Figure 6: Visualising the attentional head for the max retrieval task with (below) and without (above) adaptive temperature applied, for the same batch and parameters as in Figure 2. Note the increased sharpness in the coefficients, especially as the amount of items increases.
These results are further supplemented by a qualitative comparison of the softmax coefficients before and after applying the temperature adaptation. As can be seen in Figure 6, our proposed adaptive temperature adaptation indeed leads to sharper coefficients out-of-distribution and higher attention being directed to the desired item, even in situations where it did not receive the largest logit.
We have now successfully validated the predictions of our theory in a controlled environment. What about a more challenging benchmark with a baseline model comprising many attentional heads?
4.2 CLRS-Text
Figure 7: Resampling test results on CLRS-Text of variants of Gemma 2B, fine-tuned with and without adaptive temperature applied, on various problem sizes. Both Gemma 2B variants were explicitly trained on CLRS-Text tasks—the training set sizes are denoted by red dots—and are evaluated zero-shot. Note that we limit our sample length to $2,048$ tokens, and only show performance metrics for sizes where the answer fits in this constraint.
In this benchmark, we follow the protocol established by [MMI+24] and fine-tune Gemma 2B models [GMH+24] on the thirty algorithmic execution tasks in CLRS-Text, plotting their performance profiles in- and out-of-distribution at various problem sizes.
While it may be tempting to directly re-apply our learned adaptive temperature function from Figure 5 solely at inference time—the same way we did in the max retrieval experiments—this approach does not empirically work well in the CLRS-Text regime. This is due to the fact that CLRS-Text inputs are often textual representations of floating-point numbers and therefore individual numbers often span multiple tokens. It is therefore insufficient and inappropriate to aim for entropy levels where all the focus would be on one token only, as was desirable in the max retrieval task.
One follow-up on this could be to perform exactly the same polynomial fit exercise leading up to Figure 5, only this time focussing on “optimal” values of temperature for Gemma’s attentional heads. However, in this regime, we argue this exercise is substantially less trivial to do—as we are now dealing with a system spanning many attentional heads across many layers, it is not easy to even discover relevant attentional heads’ behaviours, and even less so to ascertain that the model’s robustness depends on those specific heads in those ways. As briefly discussed before, any such individual endeavour typically leads to a brand-new research project in mechanistic interpretability, and we do not find this to be in-scope of our paper.
That being said, there is an alternate route to make the Gemma model still benefit from our adaptive temperature module exactly as-is; it just has to directly learn how to leverage it. As such, in our CLRS-Text ablation we apply adaptive temperature both during fine-tuning and at inference time.
These final comparative results may be found in Figure 7, and they demonstrate a significant advantage of the adaptive temperature-backed model on nearly all of the thirty algorithms study. This indicates that, even in a complex system with many interactions between attentional heads, it is possible to extract benefits from the simple idea of dynamically adapting the temperature—and we hope our result paves the way for more involved future investigation of such approaches.
5 Conclusions
“Energy continuously flows from being concentrated
To becoming dispersed, spread out, wasted and useless.”
—The 2nd Law: Unsustainable, by Muse
In this paper, we have provided extensive theoretical and empirical evidence that the softmax—a key function in the design of modern frontier architectures—is fundamentally unable to sustain robust reasoning behaviours across all possible inputs, as its output coefficients are necessarily dispersing provided sufficient input elements.
Beyond illustrating and proving these dispersion effects, we also attempted to use our theoretical framework to propose an adaptive temperature approach that is able—at least to a certain extent—to hold the dispersion effect at bay. It is our opinion that the favourable results we observe with adaptive temperature warrant further investigation, and indicate that such adaptive layers are a strategy worth dedicating further attention to in future work.
We conclude by remarking, once again, that adaptive temperature is merely an ad-hoc method and it does not escape the conclusions of our theory! The key takeaway of our paper is not the adaptive temperature proposal; it is the fact that we find it worthwhile to more seriously invest in research of hybrid architectures that will not fully rely on the softmax function, at least within the confines of the assumptions of our theory. To name a few possibilities:
  • Any kind of unnormalised attention, such as linear [Sch92] or sigmoid attention [RDD+24] does not have the dispersion issues presented here. That being said, it becomes substantially harder to meaningfully rank items using them, see e.g. the GATv2 paper [BAY22].
  • Similarly, forcing the attention to be hard or local would also escape the confines of our theory. We already briefly discussed the challenges of learning with hard attention—local attention provides a very interesting alternative, but it must be stressed that “out-of-distribution” behaviours for certain heads may appear even at highly “local” scales; OOD here refers to going outside the largest problem size the head saw at training time, not the largest context deployed at training time.
  • Lastly, our key Theorem relies on the model being built out of continuous building blocks. Inserting discontinuities in the feedforward layers—perhaps using approaches like [DvGPV24] as inspiration—would also break the assumptions of our theory, though it comes with obvious challenges to learning at scale.
While such approaches haven’t seen as much success at scale as the “vanilla” Transformer, we hope our results inspire future work into making them stable, especially for constructing reasoning systems.
Acknowledgments and Disclosure of Funding
We would like to deeply thank Daniel Johnson, the author of Penzai [Joh24]—without this library, our exploratory analyses would not be nearly as fruitful. Further, we thank Alex Matthews, Andrew Dudzik and João Araùjo for helping us with the proof of one of our propositions, and Arthur Conmy, Neel Nanda and Csaba Szepesvári for reviewing the paper prior to submission and having a plethora of highly useful comments and references. Lastly, we show deep appreciation to Alex Vitvitskyi, Olga Kozlova and Larisa Markeeva, for their endless engineering support with CLRS-Text.
Appendix A Experimental details for the maximum entry retrieval task
As briefly described in the main paper, we leverage the max retrieval task over a single attention head as a way to empirically validate our theory, as well as assess the benefits of adaptive temperature in a controlled setting. In this section, we describe the various aspects of our experimental setup, for the purposes of clarity and reproducibility.
A.1 Motivation
We deliberately focus on a single attention head environment and a simple selection function (max) to remove any confounders from our observations.
Since we are using exactly one attention head, whatever coefficients it outputs can be directly related to the network’s belief in which items are most important for the downstream prediction. This allows us to, e.g., correlate the coefficients with the ground-truth magnitude of the items.
The fact that we are looking for the maximal element’s property means we are not requiring any complicated behaviour from the coefficients: when our target task is to approximate max, the softmax coefficients need to approximate argmax—which is exactly what they are designed to be a smooth approximation for. As such, this choice of target task exhibits high algorithmic alignment [XLZ+20].
A.2 Data generation
Let $n$ be the number of items in the set that we wish to classify. For each item, $1\leq i\leq n$, we need to define a priority value, which is used to select the maximal entry. We sample these values from a uniform distribution; $\rho_{i}\sim\mathcal{U}(0,1)$.
We would also wish our task to be a classification rather than regression task, in order to leverage a more robust accuracy metric. As such, let $C$ be the desired number of classes. We can now attach to each item a class, $\kappa_{i}\sim\mathcal{U}\{1,\dots,C\}$, sampled uniformly at random. In all our experiments, $C=10$.
Then, for each input item, $1\leq i\leq n$, we consider its features to be $\mathbf{x}_{i}\in\mathbb{R}^{C+1}$ to be defined as $\mathbf{x}_{i}=\rho_{i}\|\mathrm{onehot}(\kappa_{i},C)$, i.e. the concatenation of these two sampled pieces of data where $\kappa_{i}$ is represented as a one-hot vector.
Lastly, since we will leverage dot-product attention, we also need a query vector. In this particular task, the query is irrelevant, and we initialise it to a random uniformly-sampled value, $q\sim\mathcal{U}(0,1)$.
Our task is to predict, given $\{\mathbf{x}_{i}\}_{1\leq i\leq n}$ and $q$, the class of the maximal item, i.e., $\kappa_{\operatornamewithlimits{argmax}_{i}\rho_{i}}$.
A.3 Neural network architecture
The neural network model is designed to be a simple set aggregation model (in the style of Deep Sets [ZKR+17]), with a single-head dot product attention as the aggregation function.
Its equations can be summarised as follows:
$\mathbf{h}_{i}$ $=\psi_{x}(\mathbf{x}_{i})$ (5)
$\mathbf{q}$ $=\psi_{q}(q)$ (6)
$e_{i}$ $=(\mathbf{Q}\mathbf{q})^{\top}(\mathbf{K}\mathbf{h}_{i})$ (7)
$\alpha_{i}$ $=\frac{\exp(e_{i}/\theta)}{\sum_{1\leq j\leq n}\exp(e_{j}/\theta)}$ (8)
$\mathbf{z}$ $=\sum_{1\leq i\leq n}\alpha_{i}\mathbf{V}\mathbf{h}_{i}$ (9)
$\mathbf{y}$ $=\phi(\mathbf{z})$ (10)
Equations 2–3 prepare the embeddings of the items and query, using two-layer MLPs $\psi_{x}$ and $\psi_{q}$ using the GeLU activation function [HG16] and an embedding size of $128$ dimensions. Then, a single-head dot-product attention (with query, key and value matrices $\mathbf{Q}$, $\mathbf{K}$ and $\mathbf{V}$) is executed in equations 4–6. Lastly, the output class logits are predicted from the attended vector using a two-layer GeLU MLP, $\phi$. Each component is a two-layer MLP to ensure it has universal approximation properties.
A concise implementation of our network using JAX [BFH+18] and Flax [HLO+24] is as follows:
{Verbatim}
[commandchars=
{}] import jax.numpy as jnp from flax import linen as nn from typing import Callable
class Model(nn.Module): n_classes: int = 10 n_feats: int = 128 activation: Callable = nn.gelu
@nn.compact def __call__(self, x, q): x = nn.Dense(features=self.n_feats)(x) x = self.activation(x) x = nn.Dense(features=self.n_feats)(x) x = self.activation(x) q = nn.Dense(features=self.n_feats)(q) q = self.activation(q) q = nn.Dense(features=self.n_feats)(q) x = nn.MultiHeadDotProductAttention( num_heads=1, qkv_features=self.n_feats)( inputs_q=q, inputs_kv=x) x = nn.Dense(features=self.n_feats)(jnp.squeeze(x, -2)) x = self.activation(x) x = nn.Dense(features=self.n_classes)(x) return x
A.4 Experimental hyperparameters
We train our model for $100,000$ gradient steps using the Adam SGD optimiser [KB15] with initial learning rate of $\eta=0.001$. At each step, we present to the model a batch of $128$ input sets. All sets within a batch have the same size, sampled uniformly from $n\sim\mathcal{U}\{5,\dots,16\}$. The model is trained using a cross-entropy loss, along with $L_{2}$ regularisation with hyperparameter $\lambda=0.001$.
The mixed-size training is a known tactic, designed to better prepare the model for distribution shifts on larger sets at inference time. Similarly, the weight decay follows the recommendation in Proposition 3.1, as an attempt to mitigate overfitting out-of-distribution as a byproduct of sharpening the softmax coefficients.
Both methods prove to be effective in deriving a stable baseline model.
Appendix B Proof of Proposition 3.1
Proposition 3.1 (Sharpness in Transformers necessitates large weights). Let $\mathbf{e}^{(n)}\in\mathbb{R}^{n}$ be a collection of $n$ logits, computed using a dot product attention mechanism; i.e. $e^{(n)}_{k}=\left\langle\mathbf{Q}\mathbf{y},\mathbf{K}\mathbf{x}_{k}\right\rangle$, where $\mathbf{y}\in\mathbb{R}^{m}$ is a query vector and $\mathbf{Q},\mathbf{K}\in\mathbb{R}^{m^{\prime}\times m}$ are parameters. Let $\delta=\max\limits_{1\leq i\leq n}e^{(n)}_{i}-\min\limits_{1\leq j\leq n}e^{(n)}_{j}$ be their maximum difference. Then $\delta$ is upper bounded as:
$\delta\leq 2\sigma_{\mathrm{max}}^{(Q)}\sigma_{\mathrm{max}}^{(K)}\|\mathbf{y}\|\max_{1\leq i\leq n}\|\mathbf{x}_{i}\|$
where $\sigma_{\mathrm{max}}^{(Q)},\sigma_{\mathrm{max}}^{(K)}\in\mathbb{R}$ are the largest singular values of $\mathbf{Q}$ and $\mathbf{K}$. That is, the sharpness of the softmax in Transformers depends on the norm of its parameters.
Proof.
We start by showing that the largest singular values of $\mathbf{Q}$ and $\mathbf{K}$ determine the maximum stretch due to that matrix acting on $\mathbf{x}\in\mathbb{R}^{m}$. More precisely, we wish to show:
$\|\mathbf{Q}\mathbf{x}\|\leq\sigma^{(Q)}_{\mathrm{max}}\|\mathbf{x}\|\qquad\qquad\|\mathbf{K}\mathbf{x}\|\leq\sigma^{(K)}_{\mathrm{max}}\|\mathbf{x}\|$
[…]
Consider the singular value decomposition (SVD) $\mathbf{Q}=\mathbf{U}\boldsymbol{\Sigma}\mathbf{V}^{\top}$, where $\boldsymbol{\Sigma}$ is a rectangular diagonal matrix of singular values $\sigma^{(Q)}_{i}\in\mathbb{R}$. As $\mathbf{U}$ and $\mathbf{V}$ are orthogonal, $\|\mathbf{U}\mathbf{x}\|=\|\mathbf{V}\mathbf{x}\|=\|\mathbf{x}\|$. Therefore, $\|\mathbf{Q}\mathbf{x}\|=\|\mathbf{U}\boldsymbol{\Sigma}\mathbf{V}^{\top}\mathbf{x}\|=\|\boldsymbol{\Sigma}\mathbf{v}\|$, where $\mathbf{v}=\mathbf{V}^{\top}\mathbf{x}$, meaning that $\|\mathbf{v}\|=\|\mathbf{x}\|$. Then we derive:
$\|\boldsymbol{\Sigma}\mathbf{v}\|=\|\mathbf{Q}\mathbf{x}\|=\sqrt{\sum_{i}\left(\sigma^{(Q)}_{i}v_{i}\right)^{2}}\leq\sigma^{(Q)}_{\mathrm{max}}\sqrt{\sum_{i}v_{i}^{2}}=\sigma^{(Q)}_{\mathrm{max}}\|\mathbf{x}\|$
We now note that
$e_{k}^{(n)}=\left\langle\mathbf{Q}\mathbf{y},\mathbf{K}\mathbf{x}_{k}\right\rangle=\|\mathbf{Q}\mathbf{y}\|\|\mathbf{K}\mathbf{x}_{k}\|\cos\theta$
with $\theta$ the angle between the arguments of the inner product. We can now bound $e_{k}^{(n)}$ from above:
$e_{k}^{(n)}\leq\|\mathbf{Q}\mathbf{y}\|\|\mathbf{K}\mathbf{x}_{k}\|\leq\sigma_{\mathrm{max}}^{(Q)}\sigma_{\mathrm{max}}^{(K)}\|\mathbf{y}\|\|\mathbf{x}_{k}\|$
with $\sigma_{\mathrm{max}}^{(Q)},\sigma_{\mathrm{max}}^{(K)}$ being the maximum singular value of $\mathbf{Q}$ and $\mathbf{K}$, respectively, and where the last step comes from the inequality shown above. Similarly, we obtain a lower bound, yielding:
$-\sigma_{\mathrm{max}}^{(Q)}\sigma_{\mathrm{max}}^{(K)}\|\mathbf{y}\|\|\mathbf{x}_{k}\|\leq e_{k}^{(n)}\leq\sigma_{\mathrm{max}}^{(Q)}\sigma_{\mathrm{max}}^{(K)}\|\mathbf{y}\|\|\mathbf{x}_{k}\|$
This gives us the desired upper bound for $\delta$:
$\delta$ $=\max_{1\leq i\leq n}e^{(n)}_{i}-\min_{1\leq j\leq n}e^{(n)}_{j}$
$\leq\max_{1\leq i\leq n}\sigma_{\mathrm{max}}^{(Q)}\sigma_{\mathrm{max}}^{(K)}\|\mathbf{y}\|\|\mathbf{x}_{i}\|-\min_{1\leq j\leq n}-\sigma_{\mathrm{max}}^{(Q)}\sigma_{\mathrm{max}}^{(K)}\|\mathbf{y}\|\|\mathbf{x}_{j}\|$
$=\sigma_{\mathrm{max}}^{(Q)}\sigma_{\mathrm{max}}^{(K)}\|\mathbf{y}\|\max_{1\leq i\leq n}\|\mathbf{x}_{i}\|+\sigma_{\mathrm{max}}^{(Q)}\sigma_{\mathrm{max}}^{(K)}\|\mathbf{y}\|\max_{1\leq j\leq n}\|\mathbf{x}_{j}\|$
$=2\sigma_{\mathrm{max}}^{(Q)}\sigma_{\mathrm{max}}^{(K)}\|\mathbf{y}\|\max_{1\leq i\leq n}\|\mathbf{x}_{i}\|$
completing the proof. ∎
Appendix C Proof of Proposition 3.2
Proposition 3.2 (Decreasing temperature decreases entropy). Let $\mathbf{e}^{(n)}\in\mathbb{R}^{n}$ be a collection of $n$ logits. Consider the Boltzmann distribution over these $n$ items, $p_{i}\propto\exp(-\beta e_{i}^{(n)})$ for $\beta\in\mathbb{R}$, and let $H=-\sum_{i}p_{i}\log p_{i}$ be its Shannon entropy. Then, as $\beta$’s magnitude increases, $H$ must monotonically decrease. Thus, since $\beta\propto\frac{1}{\theta}$ where $\theta$ is the temperature in $\mathtt{softmax}_{\theta}$, decreasing the temperature must monotonically decrease the entropy.
Proof.
We start by briefly acknowledging the extremal values of $\beta$: at $\beta=0$ (i.e., $\theta\rightarrow\infty$), all logits are weighed equally, hence $p_{i}=\mathcal{U}(n)$ are uniform, and entropy is maximised. Similarly, at $\beta\rightarrow\pm\infty$ (i.e., $\theta=0$), either the minimum or the maximum logit is given a probability of $1$, leading to a distribution with minimal (zero) entropy.
Now, consider the partition function $Z=\sum_{i}\exp(-\beta e_{i}^{(n)})$, such that $p_{i}=\frac{\exp(-\beta e_{i}^{(n)})}{Z}$. We will take derivatives of $\log Z$ with respect to $\beta$. Starting with the first derivative:
$\frac{d}{d\beta}\log Z=\frac{1}{Z}\sum_{i}-e_{i}^{(n)}\exp(-\beta e_{i}^{(n)})=-\sum_{i}e_{i}^{(n)}p_{i}=-\mathbb{E}_{i\sim p_{i}}(e^{(n)}_{i})$
we recover the expected logit value sampled under the distribution. Now we differentiate again:
$\frac{d^{2}}{d\beta^{2}}\log Z$ $=-\frac{d}{d\beta}\sum_{i}e_{i}^{(n)}p_{i}$
$=-\sum_{i}e_{i}^{(n)}\frac{d}{d\beta}\frac{\exp(-\beta e_{i}^{(n)})}{Z}$
$=-\sum_{i}e_{i}^{(n)}\frac{-e_{i}^{(n)}\exp(-\beta e_{i}^{(n)})Z-\exp(-\beta e_{i}^{(n)})\sum_{j}-e_{j}^{(n)}\exp(-\beta e_{j}^{(n)})}{Z^{2}}$
$=\sum_{i}(e_{i}^{(n)})^{2}\frac{\exp(-\beta e_{i}^{(n)})}{Z}-\sum_{j}e_{j}^{(n)}\frac{\exp(-\beta e_{j}^{(n)})}{Z}\frac{\sum_{k}e_{k}^{(n)}\exp(-\beta e_{k}^{(n)})}{Z}$
$=\sum_{i}(e_{i}^{(n)})^{2}p_{i}-\sum_{j}e_{j}^{(n)}p_{j}\sum_{k}e_{k}^{(n)}p_{k}$
$=\mathbb{E}_{i\sim p_{i}}((e_{i}^{(n)})^{2})-\mathbb{E}_{i\sim p_{i}}(e_{i}^{(n)})^{2}=\mathrm{Var}_{i\sim p_{i}}(e_{i}^{(n)})$
and we recover the variance of the expected logit value.
Now we turn our attention to the entropy formula:
$H=-\sum_{i}p_{i}\log p_{i}$ $=-\sum_{i}p_{i}(\log\exp(-\beta e_{i}^{(n)})-\log Z)$
$=\sum_{i}p_{i}\log Z-\sum_{j}-\beta e_{j}^{(n)}p_{j}$
$=\log Z+\beta\mathbb{E}_{i\sim p_{i}}(e_{i}^{(n)})=\log Z-\beta\frac{d}{d\beta}\log Z$
To check the monotonicity of $H$ as $\beta$ varies, we now take the derivative of this expression w.r.t. $\beta$:
$\frac{dH}{d\beta}=\frac{d}{d\beta}\log Z-\frac{d}{d\beta}\log Z-\beta\frac{d^{2}}{d\beta^{2}}\log Z=-\beta\frac{d^{2}}{d\beta^{2}}\log Z=-\beta\mathrm{Var}_{i\sim p_{i}}(e_{i}^{(n)})$
Since variance can never be negative, we find that $\frac{dH}{d\beta}\leq 0$ when $\beta\geq 0$, and $-\frac{dH}{d\beta}\leq 0$ when $\beta\leq 0$. As such, as the magnitude $|\beta|$ grows, the value of $H$ must monotonically decrease. ∎